find the lowest common denominator of x^2+12x+36 and x^2+3x-18
Answer by jorel1380(3719)

x^2+12x+36 and x^2+3x-18
x^2+12x+36=(x+6)2
x^2+3x-18=(x+6)(x-3)
LCD=(x+6)..